# Asymptotic independence in higher dimensions and its implications on risk management

### v2

@@ Abstract
In the study of extremes, the presence of asymptotic independence signifies that extreme events across multiple variables are probably less likely to occur together. Although well-understood in a bivariate context, the concept remains relatively unexplored when addressing the nuances of joint occurrence of extremes in higher dimensions. In this paper, we propose a notion of mutual asymptotic independence to capture the behavior of joint extremes in dimensions larger than two and contrast it with the classical notion of (pairwise) asymptotic independence. Additionally, we define $k$-wise asymptotic independence, which captures the tail dependence between pairwise and mutual asymptotic independence. The concepts are compared using examples of Archimedean, Gaussian, and Marshall-Olkin copulas among others. Notably, for the popular Gaussian copula, we provide explicit conditions on the correlation matrix for mutual asymptotic independence and $k$-wise asymptotic independence to hold; moreover, we are able to compute exact tail orders for various tail events. Beside that, we compare and discuss the implications of these new notions of asymptotic independence on assessing the risk of complex systems under distributional ambiguity.
62H05,
62H20,
62G32,
62G32, 62P05,
asymptotic independence,
copula models,
Gaussian copula,
multivariate extremes,
@@ 1 Introduction
In many multivariate models, we observe that the likelihood of joint occurrence of extreme values in two or more variables is negligible in comparison to the occurrence of an extreme value in one variable. In this context, the notion of asymptotic independence looms large in the study of joint extreme values in probability distributions, although mostly restricted to the bivariate set-up. A random vector $(Z_{1},Z_{2})\in\mathbb{R}^{2}$ with identically distributed marginals is asymptotically (right-tail/upper-tail) independent if
@@
Often called Sibuya’s condition, (1.1) was exhibited by Sibuya [38] for bivariate normal random vectors with any correlation $\rho<1$. Such a limit behavior has also been found to hold for bivariate distributions with arbitrary choice of marginals possessing a variety of dependence structures, including, Frank copula, Ali-Mikhail-Haq copula, Gaussian copula, Farlie-Gumbel-Morgenstern copula, and, more; see [28, 29, 5, 21]. It is widely believed that the presence of asymptotic independence hinders the computation of joint tail probabilities, and has led to a variety of techniques for modeling and estimating rare tail probabilities when such a property is present; see [28, 37, 36, 30, 9, 7]. Nevertheless, for random vectors in dimensions higher than two, limited expositions are available, and multivariate asymptotic independence is often understood to be (1.1) holding for all pairs of variables, which we call pairwise asymptotic independence.
@@
where $11/2\leq\eta<$ 1 and $\ell$ is slowly varying at infinity (i.e., $\ell(tz)/\ell(z)\to 1$ as $t\to\infty$, $\forall\,z>0$), then $\eta$ represents this coefficient of tail dependence. According to Ledford and Tawn [28], (i) $\eta=1/2$ and $\ell(t)\geq 1$ signifies near independence, (ii) $\eta=1$ and $\ell(t)\nrightarrow 0$ as $t\to\infty$ signifies upper tail dependence, and finally, (iii) either $11/2<\eta<$ 1, or $\eta=1$ and $\ell(t)\to 0$ as $t\to\infty$ signifies positive association.
@@
A copula $C:[0,1]^{d}\to[0,1]$ is a multivariate distribution function with identical uniform $[0,1]$ margins. From Sklar’s Theorem [39, 34, 14] we know that for any $d$-dimensional random vector $\boldsymbol{Z}=(Z_{1},\ldots,Z_{d})$ with distribution function $F$ and marginal distributions $F_{1},\ldots,F_{d}$ there exists a copula $C:[0,1]^{d}\to[0,1]$ such that
@@
for $(z_{1},\ldots,z_{d})\in\mathbb{R}^{d}$, and if the margins are continuous, the copula is uniquely given by
@@
is the generalized inverse of $F_{j}$ for $j=1,\ldots,d$. In this paper, we are particularly concerned with the probability of joint extremes where the survival copula $\widehat{C}:[0,1]^{d}\to[0,1]$, which is also a copula, plays an important role. The survival copula $\widehat{C}$ satisfies
@@
Note that for independence of multiple random variables, it is well-known that “pairwise independence” for all pairs of random variables is not equivalent to their “mutual independence” (cf. [22, Chapter 2]). In the same vein, we propose here the concepts of pairwise asymptotic independence in Section 2 and mutual asymptotic independence in Section 3. With the new notion of mutual asymptotic independence, we explore the ideas of “near independence” and “intermediate upper tail dependence” through all subsequent dimensions $2,3,\ldots,d$ going beyond just the $d$-dimensional characterization as given in (1.2). For models that lie between pairwise and mutually asymptotically independent models, we introduce the concept of $k$-wise asymptotic independence for $k\in\{2,\ldots,d\}$ in Section 4. In particular, we investigate and compare the various notions of asymptotic independence and illustrate them using popular copula models. Moreover, we obtain the following three key results for the popular Gaussian copula, which have broader theoretical and practical implications:
@@
    a derivation of the correct tail orders for various tail events, and,
@@
Besides the Gaussian copula, we also provide examples to exhibit the breadth of asymptotic (in)dependence behavior using the Archimedean copula family. We apply the new notions of asymptotic independence in Section 5 to show its implications on assessing the risk of complex systems under distributional ambiguity. The different notions of asymptotic independence influence risk contagion in financial systems differently and hence may lead to an underestimation or overestimation of risk if applied improperly. In particular, we exhibit this phenomenon using two pertinent conditional risk measures, namely, conditional tail probability and Contagion value-at-risk or CoVaR in dimensions $d>2$. Finally, in Section 6, we conclude with some broader implications of interpreting asymptotic independence in this new light. All proofs for the results presented in this paper are provided in Appendix A.
@@
We denote by $\mathbb{I}_{d}=\{1,\ldots,d\}$ an index set with $d$ elements and the cardinality of a set $S\subseteq\mathbb{I}_{d}$ is denoted by $|S|$. For a random vector $\boldsymbol{Z}=(Z_{1},\ldots,Z_{d})$, we write $\boldsymbol{Z}\sim F$ if $\boldsymbol{Z}$ has distribution function $F$; moreover, we understand that marginally $Z_{j}\sim F_{j}$ for $j\in\mathbb{I}_{d}$. For any non-empty sets $S\subseteq\mathbb{I}_{d}$, the copula and survival copula of the corresponding $|S|$-dimensional marginal are denoted by $C_{S}$ and $\widehat{C}_{S}$, respectively. Moreover, if $d=1$ we have $C_{S}(u)=\widehat{C}_{S}(u)=u$ for $0\leq u\leq 1$. For a given vector $\boldsymbol{z}\in\mathbb{R}^{d}$ and $S\subseteq\mathbb{I}_{d}$, we denote by $\boldsymbol{z}^{\top}$ the transpose of $\boldsymbol{z}$ and by $\boldsymbol{z}_{S}\in\mathbb{R}^{|S|}$ the vector obtained by deleting the components of $\boldsymbol{z}$ in $\mathbb{I}_{d}\backslash S$. Similarly, for non-empty $S\subseteq\mathbb{I}_{d}$, $\Sigma_{S}$ denotes the appropriate sub-matrix of a given matrix $\Sigma\in\mathbb{R}^{d\times d}$ after removing all rows and columns with indices in $\mathbb{I}_{d}\setminus S$. Furthermore, $\boldsymbol{0}_{d}=(0,\ldots,0)^{\top}$ and $\boldsymbol{1}_{d}=(1,\ldots,1)^{\top}$ are vectors in $\mathbb{R}^{d}$, and $\boldsymbol{I}_{d}$ is the identity matrix in $\mathbb{R}^{d\times d}$; subscripts are dropped when evident from the context. Vector operations are understood component-wise, e.g., for vectors $\boldsymbol{z}=(z_{1},\ldots,z_{d})$ and $\boldsymbol{y}=(y_{1},\ldots,y_{d})$, $\boldsymbol{z}\leq\boldsymbol{y}$ means $z_{j}\leq y_{j}$, $\forall j\in\mathbb{I}_{d}$. For functions $f,g:(0,\infty)\to(0,\infty)$, we write $f(u)\sim g(u)$ as $u\downarrow 0$ if $\lim_{u\downarrow 0}f(u)/g(u)=1.$ Moreover, a function $\ell:(0,\infty)\to(0,\infty)$ is called a slowly varying at 00, if $\lim_{u\downarrow 0}\ell(uz)/\ell(u)=1,\forall z>0$.
@@ 2 Pairwise asymptotic independence
Note that the definition in (1.1) can be easily generalized to distributions with potentially unequal margins; any random vector $(Z_{1},Z_{2})$ with continuous marginals $Z_{j}\sim F_{j},$ $j=1,2$ is asymptotically independent if
@@
An interesting feature of this definition of asymptotic independence is that it is based on tail sets tethered along the main diagonal $(t,t)$ (in (1.1)) or $(1-u,1-u)$ (in (2.1)). It is easy to check that (2.1) is equivalent to
@@
for some $a,b>0$ (cf. [2, Theorem 2]). Curiously, an equivalent result for the distribution function of a bivariate random vector with survival function $\widehat{C}$ does not hold: even if (1.1) holds it does not necessarily hold for diagonals of the form $(at,bt)$ for any $a,b>0$; see Das and Fasen-Hartmann [11, Proposition 3.9] for an example with normally distributed marginals $(Z_{1},Z_{2})$ where $\mathbb{P}(Z_{1}>at,Z_{2}>t)=O(\mathbb{P}(Z_{2}>t)),$ as $t\to\infty$.
@@
A random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ with survival copula $\widehat{C}$ is pairwise asymptotically independent if $\forall\,j,\ell\in\mathbb{I}_{d},j\neq\ell,$
$\widehat{C}_{j,\ell}(u,u)=o(u),\quad\quad u\downarrow 0.$ (2.2)
@@
Although we assumed continuous marginals for stating the condition in (2.1), in this paper, we will be primarily concerned with the asymptotic behavior of the survival copula, hence the marginal behavior and its continuity will not be relevant henceforth.
Remark 2.3.
@@
$\lim_{u\downarrow 0}\frac{\widehat{C}(u,\ldots,u)}{u}=K\in(0,\infty).$ (2.3)
@@ 2.1 Examples
$\widehat{C}_{j,\ell}^{\text{ind}}{(u_{1},u_{2})=u_{1}u_{2},\quad 0\leq u_{1},u_{2}\leq 1}.$
@@
In particular, for any distinct $j,\ell\in\mathbb{I}_{d}$, we have
$\widehat{C}^{\mathrm{MO}(\Lambda)}_{\{i,j\}}(u,u)=u^{\eta^{*}}$
@@
$\eta^{*}$ $=\sum_{S\subseteq\mathbb{I}_{d}\atop 1\in S,2\notin S}\eta_{1}^{S}+\sum_{S\subseteq\mathbb{I}_{d}\atop 1\notin S,2\in S}\eta_{2}^{S}+\sum_{S\subseteq\mathbb{I}_{d}\atop 1,2\in S}\max\{\eta_{1}^{S},\eta_{2}^{S}\}.$ (2.7)
Proposition 2.6.
The generalized Marshall-Olkin dependence as given by the survival copula $\widehat{C}^{\mathrm{MO}(\Lambda)}$ in (2.5) possesses pairwise asymptotic independence for any choice of $\Lambda$.
@@
    Equal parameter for all sets: Here, $\lambda_{S}=\lambda$ for all non-empty $S\subseteq\mathbb{I}_{d}$ where $\lambda>0$ and we denote the survival copula by $\widehat{C}^{\mathrm{MO}^{=}}$. We can check from (2.6) that the value of $\lambda$ is irrelevant here. Clearly in this case $\eta_{j}^{S}=1/2^{d-1},$ for all $j\in S$ and non-empty $S\subset\mathbb{I}_{d}$. Hence we can compute the value of $\eta^{*}$ defined in (2.7) as
    $\eta^{*}$ $=\sum_{S\subseteq\mathbb{I}_{d}\atop 1\in S}\frac{1}{2^{d-1}}+\sum_{S\subseteq\mathbb{I}_{d}\atop 1\notin S,2\in S}\frac{1}{2^{d-1}}=\frac{2^{d-1}+2^{d-2}}{2^{d-1}}=\frac{3}{2}.$
    Therefore, for all $j,\ell\in S$ with $j\not=\ell$,
    $\widehat{C}^{\mathrm{MO}^{=}}_{j,\ell}(u,u)$ $=u^{3/2},\quad 0\leq u\leq 1.$
@@
    Parameters proportional to the cardinality of the sets: Here, $\lambda_{S}=|S|\lambda$ for all non-empty $S\subseteq\mathbb{I}_{d}$ where $\lambda>0$ and we denote the survival copula by $\widehat{C}^{\mathrm{MO}^{\infty}}$. As well the value of $\lambda$ is irrelevant and for all $j\in S$ and non-empty subset $S\subset\mathbb{I}_{d}$ we have
    $\eta_{j}^{S}=\frac{|S|}{(d+1)2^{d-2}}.$
    We compute again the value of $\eta^{*}$ defined in (2.7) as
    $\eta^{*}$ $=\sum_{S\subseteq\mathbb{I}_{d}\atop 1\in S}\frac{|S|}{(d+1)2^{d-2}}+\sum_{S\subseteq\mathbb{I}_{d}\atop 1\notin S,2\in S}\frac{|S|}{(d+1)2^{d-2}}$
@@
    Therefore, for all $j,\ell\in S$ with $j\not=\ell$,
    $\widehat{C}^{\mathrm{MO}^{\infty}}_{j,\ell}(u,u)$ $=u^{1+d/(2(d+1))},\quad 0\leq u\leq 1.$
@@
A useful family of copula models for multivariate distributions is the Archimedean copulas ([25, 4]). A $d$-dimensional copula $C^{\phi}$ is Archimedean if
@@
for $0\leq u_{1},\ldots,u_{d}\leq 1$, where the generator function $\phi:[0,1]\to[0,\infty]$ is convex, decreasing, with $\phi(1)=0$ and $\phi^{{\leftarrow}}(y)=\inf\{u\in[0,1]:\phi(u)\leq y\}$ for $y\in(0,\infty)$. Necessary and sufficient conditions on the function $\phi$ such that $C^{\phi}$ in (2.8) is a copula is given in [33]; note that the survival copula $\widehat{C}^{\phi}$ of an Archimedean copula $C^{\phi}$ is in general not Archimedean. A popular choice of $\phi$ is the Laplace transform of any positive random variable.
Tail dependence in such copulas has been studied in [4, 23], and sufficient conditions to obtain pairwise independence exist. Suppose the random vector $\boldsymbol{Z}$ having an Archimedean copula $C^{\phi}$ has a generator $\phi$ satisfying
@@
exists and is larger than $1$ then we have asymptotic upper tail dependence, i.e, $\widehat{C}^{\phi}$ satisfies (2.3). We observe from Charpentier and Segers [4, Table 1] that for many Archimedean copulas, we have $\theta_{1}=1$ and thus they are pairwise asymptotically independent; this includes Frank copula, Clayton copula, Ali-Mikhail-Haq copula and so on; see also Nelsen [34, Table 4.1] for further details.
@@
The Gaussian dependence structure is perhaps the most popular one used in practice. Let $\Phi_{\Sigma}$ denote the distribution function of a $d$-variate normal distribution with all marginal means zero, variances one and positive-definite correlation matrix $\Sigma\in\mathbb{R}^{d\times d}$, and $\Phi$ denote a standard normal distribution function. Then for $10<u_{1},\ldots,u_{d}<$ 1,
@@
denotes the Gaussian copula with correlation matrix $\Sigma$.
Pairwise asymptotic independence has been well-known for the bivariate normal distribution, as well as the bivariate Gaussian copula ([38, 28]). But not surprisingly, it holds as well for $d\geq 2$ as we summarize next.
Proposition 2.9.
If the dependence structure of a random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ is given by a Gaussian copula $C^{\Sigma}$ with positive-definite correlation matrix $\Sigma$, then $\boldsymbol{Z}$ exhibits pairwise asymptotic independence.
In fact, it is possible to find the exact tail order for the Gaussian survival copula for any $S\subseteq\mathbb{I_{d}}$ with $|S|\geq 2$, the precise result is given in Section 3.2.2.

### v3

@@ Abstract
In the study of extremes, the presence of asymptotic independence signifies that extreme events across multiple variables are probably less likely to occur together. Although well-understood in a bivariate context, the concept remains relatively unexplored when addressing the nuances of the joint occurrence of extremes in higher dimensions. In this paper, we propose a notion of mutual asymptotic independence to capture the behavior of joint extremes in dimensions larger than two and contrast it with the classical notion of (pairwise) asymptotic independence. Additionally, we define $k$-wise asymptotic independence, which captures the tail dependence between pairwise and mutual asymptotic independence. The concepts are compared using examples of Archimedean, Gaussian and Marshall-Olkin copulas, among others. Finally, we discuss the implications of these new notions of asymptotic independence on assessing the risk of complex systems under distributional ambiguity.
@@ 1 Introduction
In many multivariate models, we observe that the likelihood of the joint occurrence of extreme values in two or more variables is negligible in comparison to the occurrence of an extreme value in one variable. In this context, the notion of asymptotic independence looms large in the study of joint extreme values in probability distributions, although mostly restricted to the bivariate set-up. A random vector $(Z_{1},Z_{2})\in\mathbb{R}^{2}$ with identically distributed marginals is asymptotically (right-tail/upper-tail) independent if
@@
Often called Sibuya’s condition, (1.1) was exhibited by [41] for bivariate normal random vectors with any correlation $\rho<1$. Such a limit behavior has also been found to hold for bivariate distributions with arbitrary choice of marginals possessing a variety of dependence structures, including Frank copula, Ali-Mikhail-Haq copula, Gaussian copula, Farlie-Gumbel-Morgenstern copula, and more; see [29, 30, 5, 22]. It is widely believed that the presence of asymptotic independence hinders the computation of joint tail probabilities, and has led to a variety of techniques for modeling and estimating rare tail probabilities when such a property is present; see [29, 39, 38, 31, 7, 9]. Nevertheless, for random vectors in dimensions higher than two, limited expositions are available, and multivariate asymptotic independence is often understood to be (1.1) holding for all pairs of variables, which we call pairwise asymptotic independence. Such a notion of multivariate asymptotic independence possibly has its origins in the study of extremes. For instance, in Resnick [40, Chapter 5.5], a multivariate distribution is called multivariate asymptotically independent if it is in the maximum domain of attraction of a multivariate extreme value distribution with independent marginals. Additionally, it is shown that such a characterization of “multivariate asymptotic independence” is equivalent to having “pairwise asymptotic independence” (assuming identical marginals in the maximum domain of attraction of a univariate extreme value distribution); see Resnick [40, Proposition 5.27], Galambos [17, Corollary 5.3.1]. In this paper, we show that asymptotic tail independence often may have a much subtler form that goes beyond pairwise asymptotic independence (multivariate asymptotic independence).
@@
where $1/2\leq\eta<1$ and $\ell$ is slowly varying at infinity (i.e., $\ell(tz)/\ell(z)\to 1$ as $t\to\infty$, $\forall\,z>0$), then $\eta$ represents this coefficient of tail dependence. According to Ledford and Tawn [29], (i) $\eta=1/2$ and $\ell(t)\geq 1$ signifies near independence, (ii) $\eta=1$ and $\ell(t)\nrightarrow 0$ as $t\to\infty$ signifies upper tail dependence, and finally, (iii) either $1/2<\eta<1$, or $\eta=1$ and $\ell(t)\to 0$ as $t\to\infty$ signifies positive association.
@@
A copula $C:[0,1]^{d}\to[0,1]$ is a multivariate distribution function with identical uniform $[0,1]$ marginals. From Sklar’s Theorem [42, 36, 14] we know that for any $d$-dimensional random vector $\boldsymbol{Z}=(Z_{1},\ldots,Z_{d})$ with distribution function $F$ and marginal distributions $F_{1},\ldots,F_{d}$ there exists a copula $C:[0,1]^{d}\to[0,1]$ such that
@@
for $(z_{1},\ldots,z_{d})\in\mathbb{R}^{d}$, and if the marginals are continuous, the copula is uniquely given by
@@
is the generalized inverse of $F_{j}$ for $j=1,\ldots,d$. In this paper, we are particularly concerned with the probability of joint extremes where the survival copula $\widehat{C}:[0,1]^{d}\to[0,1]$, which is also a copula, plays an important role; see Durante and Sempi [14, Chapter 1], McNeil, Frey and Embrechts [34, Section 5.1.5], Nelsen [36, Section 2.6]. The survival copula $\widehat{C}$ satisfies
@@
Note that for independence of multiple random variables, it is well-known that “pairwise independence” for all pairs of random variables is not equivalent to their “mutual independence” (cf. Hogg, McKean and Craig [23, Chapter 2]). In the same vein, we propose here the concepts of pairwise asymptotic independence in Section 2 and mutual asymptotic independence in Section 3. With the new notion of mutual asymptotic independence, we explore the ideas of “near independence” and “intermediate upper tail dependence” through all subsequent dimensions $2,3,\ldots,d$ going beyond just the $d$-dimensional characterization as given in (1.2). For models that lie between pairwise and mutually asymptotically independent models, we introduce the concept of $k$-wise asymptotic independence for $k\in\{2,\ldots,d\}$ in Section 4. In particular, we investigate and compare the various notions of asymptotic independence and illustrate them using popular copula models. Moreover, we obtain the following three key results for the popular Gaussian copula, which have broader theoretical and practical implications:
@@
    a derivation of the correct tail orders, and,
@@
Besides the Gaussian copula, we also provide examples to exhibit the breadth of asymptotic (in)dependence behavior using the Archimedean copula family. We apply the new notions of asymptotic independence in Section 5 to show its implications on assessing the risk of complex systems under distributional ambiguity. The different notions of asymptotic independence influence risk contagion in financial systems differently and hence may lead to an underestimation or overestimation of risk if applied improperly. In particular, we exhibit this phenomenon using two pertinent conditional risk measures, namely, conditional tail probabilities and Contagion Value-at-Risk or Conditional Value-at-Risk (CoVaR) in dimensions $d>2$. Finally, in Section 6, we conclude with some broader implications of interpreting asymptotic independence in this new light. All proofs for the results presented in this paper are provided in the Appendix.
@@
We denote by $\mathbb{I}_{d}=\{1,\ldots,d\}$ an index set with $d$ elements and the cardinality of a set $S\subseteq\mathbb{I}_{d}$ is denoted by $|S|$. For a random vector $\boldsymbol{Z}=(Z_{1},\ldots,Z_{d})$, we write $\boldsymbol{Z}\sim F$ if $\boldsymbol{Z}$ has distribution function $F$; moreover, we understand that marginally $Z_{j}\sim F_{j}$ for $j\in\mathbb{I}_{d}$. For any non-empty sets $S\subseteq\mathbb{I}_{d}$, the copula and survival copula of the corresponding $|S|$-dimensional marginal are denoted by $C_{S}$ and $\widehat{C}_{S}$, respectively. Moreover, if $d=1$ we have $C_{S}(u)=\widehat{C}_{S}(u)=u$ for $0\leq u\leq 1$. For a given vector $\boldsymbol{z}\in\mathbb{R}^{d}$ and $S\subseteq\mathbb{I}_{d}$, we denote by $\boldsymbol{z}^{\top}$ the transpose of $\boldsymbol{z}$ and by $\boldsymbol{z}_{S}\in\mathbb{R}^{|S|}$ the vector obtained by deleting the components of $\boldsymbol{z}$ in $\S\mathbb{I}_{d}\backslash$ S. Similarly, for non-empty $S\subseteq\mathbb{I}_{d}$, $\Sigma_{S}$ denotes the appropriate sub-matrix of a given matrix $\Sigma\in\mathbb{R}^{d\times d}$ after removing all rows and columns with indices in $\mathbb{I}_{d}\setminus S$. Furthermore, $\boldsymbol{0}_{d}=(0,\ldots,0)^{\top}$ and $\boldsymbol{1}_{d}=(1,\ldots,1)^{\top}$ are vectors in $\mathbb{R}^{d}$, and $\boldsymbol{I}_{d}$ is the identity matrix in $\mathbb{R}^{d\times d}$; subscripts are dropped when evident from the context. Vector operations are understood component-wise, e.g., for vectors $\boldsymbol{z}=(z_{1},\ldots,z_{d})$ and $\boldsymbol{y}=(y_{1},\ldots,y_{d})$, $\boldsymbol{z}\leq\boldsymbol{y}$ means $z_{j}\leq y_{j}$, $\forall j\in\mathbb{I}_{d}$. For functions $f,g:(0,\infty)\to(0,\infty)$, we write $f(u)\sim g(u)$ as $u\downarrow 0$ if $\lim_{u\downarrow 0}f(u)/g(u)=1.$ Moreover, a function $\ell:(0,\infty)\to(0,\infty)$ is called a slowly varying at 0, if $\lim_{u\downarrow 0}\ell(uz)/\ell(u)=1,\forall z>0$.
@@ 2 Pairwise asymptotic independence
Note that the definition in (1.1) can be easily generalized to distributions with potentially unequal marginals; any random vector $(Z_{1},Z_{2})$ with continuous marginals $Z_{j}\sim F_{j},$ $j=1,2$ is asymptotically independent if
@@
An interesting feature of this definition of asymptotic independence is that it is based on tail sets tethered along the main diagonal $(t,t)$ (in (1.1)) or (1u,1u)(1-u,1-u) (in (2.1)). It is easy to check that (2.1) is equivalent to
@@
for some $a,b>0$ (Balkema and Nolde [2, Theorem 2]). Curiously, an equivalent result for the distribution function of a bivariate random vector does not hold: even if (1.1) holds it does not necessarily hold for diagonals of the form $(at,bt)$ for any $a,b>0$; see Das and Fasen-Hartmann [11, Proposition 3.9] for an example with normally distributed marginals $(Z_{1},Z_{2})$ where $\mathbb{P}(Z_{1}>at,Z_{2}>t)=O(\mathbb{P}(Z_{2}>t)),$ as $t\to\infty$.
@@
A random vector $\boldsymbol{Z}\in\mathbb{R}^{d}$ with continuous marginal distributions and survival copula $\widehat{C}$ is pairwise asymptotically independent if $\forall\,j,\ell\in\mathbb{I}_{d},j\neq\ell,$
$\widehat{C}_{\{j,\ell\}}(u,u)=o(u),\quad\quad u\downarrow 0.$ (2.2)
@@
  1. (a)
    Note that for discrete distributions, the copula is no longer necessarily unique, so we cannot extend this definition straightforwardly. Therefore, we assumed continuous margins to receive the uniqueness of the copula. However, in this paper, we are concerned with tail probabilities. Hence, it suffices to have the marginal distributions to be continuous above a fixed threshold (i.e., they are eventually continuous). For convenience, since we are primarily concerned with the asymptotic behavior of the survival copula, we will assume for the rest of the paper that all distributions have continuous marginal distributions.
  2. (b)
@@
    $\lim_{u\downarrow 0}\frac{\widehat{C}(u,\ldots,u)}{u}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\lambda\in(0,1)}.$ (2.3)
@@ 2.1 Examples
$\widehat{C}_{\{j,\ell\}}^{\text{ind}}{(u_{1},u_{2})=u_{1}u_{2},\quad 0\leq u_{1},u_{2}\leq 1}.$
@@
For any distinct $j,\ell\in\mathbb{I}_{d}$, we can compute that
[…]
@@
$\eta^{*}_{j\ell}$ $=\sum_{S\subseteq\mathbb{I}_{d}\atop j\in S,\ell\notin S}\eta_{j}^{S}+\sum_{S\subseteq\mathbb{I}_{d}\atop j\notin S,\ell\in S}\eta_{\ell}^{S}+\sum_{S\subseteq\mathbb{I}_{d}\atop j,\ell\in S}\max\{\eta_{j}^{S},\eta_{\ell}^{S}\}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}>1}.$ (2.7)
Clearly, since $\eta_{j\ell}^{*}>1,\forall j\neq\ell$, $\widehat{C}^{\mathrm{MO}(\Lambda)}$ possesses pairwise asymptotic independence for any choice of $\Lambda$; see as well [32, Proposition 3]. An even larger class of Marshall-Olkin copulas have been introduced in Lin and Li [32] which are also pairwise asymptotically independent.
@@
    Equal parameter for all sets: Here, $\lambda_{S}=\lambda$ for all non-empty $S\subseteq\mathbb{I}_{d}$ where $\lambda>0$ and we denote the survival copula by $\widehat{C}^{\mathrm{MO}^{=}}$. We can check from (2.6) that the value of $\lambda$ is irrelevant here. Clearly in this case $\eta_{j}^{S}=1/2^{d-1},$ for all $j\in S$ and non-empty $S\subset\mathbb{I}_{d}$. Hence we can compute the value of $\eta^{*}_{j\ell}$ defined in (2.7) as
[…]
    Therefore, for all $j,\ell\in S$ with $j\not$$\ell$,
[…]
@@
    Parameters proportional to the cardinality of the sets: Here, $\lambda_{S}=|S|\lambda$ for all non-empty $S\subseteq\mathbb{I}_{d}$ where $\lambda>0$ and we denote the survival copula by $\widehat{C}^{\mathrm{MO}^{\infty}}$. As well the value of $\lambda$ is irrelevant and for all $j\in S$ and non-empty subset $S\subset\mathbb{I}_{d}$ we have η_j^S = —S—(d+1)2d-2. We compute again the value of $\eta^{*}_{j\ell}$ defined in (2.7) as
[…]
@@
    Therefore, for all $j,\ell\in S$ with $j\not$$\ell$,
[…]
@@
A useful family of copula models for multivariate distributions is the Archimedean copulas [26, 4]. A $d$-dimensional copula $C^{\phi}$ is Archimedean if
@@
for $0\leq u_{1},\ldots,u_{d}\leq 1$, where the generator function $\phi:[0,1]\to[0,\infty]$ is convex, decreasing, with $\phi(1)=0$ and $\phi^{{\leftarrow}}(y)=\inf\{u\in[0,1]:\phi(u)\leq y\}$ for $y\in(0,\infty)$. Necessary and sufficient conditions on the function $\phi$ such that $C^{\phi}$ in (2.8) is a copula are given in [35]; note that the survival copula $\widehat{C}^{\phi}$ of an Archimedean copula $C^{\phi}$ is, in general, not Archimedean. A popular choice of $\phi$ is the Laplace transform of any positive random variable.
Tail dependence in such copulas has been studied in [4, 24], and sufficient conditions to obtain pairwise asymptotic independence exist. Suppose the random vector $\boldsymbol{Z}$ having an Archimedean copula $C^{\phi}$ has a generator $\phi$ satisfying
@@
exists and is larger than $1$ then we have asymptotic upper tail dependence, i.e, $\widehat{C}^{\phi}$ satisfies (2.3). We observe from Charpentier and Segers [4, Table 1] that for many Archimedean copulas, we have $\theta_{1}=1$ and thus they are pairwise asymptotically independent; this includes Frank copula, Clayton copula, Ali-Mikhail-Haq copula, and so on; see also Nelsen [36, Table 4.1] for further details.
@@
The Gaussian dependence structure is perhaps the most popular one used in practice. Let $\Phi_{\Sigma}$ denote the distribution function of a $d$-variate normal distribution with all marginal means zero, variances one and a positive-definite correlation matrix $\Sigma\in\mathbb{R}^{d\times d}$, and $\Phi$ denote a standard normal distribution function. Then for $0<u_{1},\ldots,u_{d}<1$,
@@
denotes the Gaussian copula with correlation matrix $\Sigma$. Pairwise asymptotic independence has been well-known for the bivariate normal distribution, as well as the bivariate Gaussian copula if the correlation is less than one [41, 29]. Hence, we may immediately conclude that for $d\geq 2$, a Gaussian copula $C^{\Sigma}$ exhibits pairwise asymptotic independence if $\Sigma$ is positive definite. In fact, it is possible to find the exact tail order for the Gaussian survival copula for any $S\subseteq\mathbb{I_{d}}$ with $|S|\geq 2$, the precise result is given in Section 3.2.2.